OpenStudy (anonymous):

log base 9 (9^(6)/27)

OpenStudy (anonymous):

\[\log_{9}(9^6/27)=\]

OpenStudy (anonymous):

I ended up with: \[\log_{9}(3^{15})\]

OpenStudy (anonymous):

\[ \log_{9} \left( \frac{9^6}{27} \right) = \log_{9} \left( \frac{9^6}{3^3} \right)\]\[= \log_{9} \left(\frac{9^6}{9^{\frac{3}{2}}}\right)\]\[= \log_{9} \left( 9^{\frac{9}{2}} \right)\]\[= \frac{9}{2}\]

OpenStudy (anonymous):

ohhhh ok I got it

OpenStudy (anonymous):

Thank exraven!

OpenStudy (anonymous):

thanks*